for f(x)=X^2+2x.find f(-1)
Found 2 solutions by jim_thompson5910, MathLover1:
Answer by jim_thompson5910(35256)

f(x) = x^2+2x

f(-1) = (-1)^2+2(-1)

f(-1) = 1+2(-1)

f(-1) = 1-2

f(-1) = -1

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by MathLover1(20849)